I have a 52' diameter cirle that i want to change into a polygon. I'm guessing it will have approximately 28 sides with each length at approximately 6' what would be the angle of each section.
Answer by Alan3354(69443)

I have a 52' diameter cirle that i want to change into a polygon. I'm guessing it will have approximately 28 sides with each length at approximately 6' what would be the angle of each section.
----------
The # of sides can't be approximate, it has to be an integer.
---
If all sides are equal length:
360/28 = 90/7 degrees =~ 12.85 degs
----
The other dimensions are not relevant.